ANALYSIS OF THE LOWEST IN-PLANE BEND AND FIRST EXCITED TORSIONAL STATE OF CH_{3}CH_{2}CN^{a}

\underline{CAROLYN S. BRAUER}, JOHN C. PEARSON, BRIAN J. DROUIN, SHANSHAN YU, JET PROPULSION LABORATORY, CALIFORNIA INSTITUTE OF TECHNOLOGY, 4800 OAK GROVE DR., PASADENA, CA 91109.

Propionitrile (CH_{3}CH_{2}CN) is observed with large column densities in a number of high-mass star-forming cores, where core temperatures exceed 200 K. It is a near-prolate (κ = 0.96) asymmetric top with appreciable dipole moment components on both the a− and b−axes (μ_{a} = 3.84 D, μ_{b} = 1.23 D).^{a} This, combined with the presence of four fundamental modes as well as four overtones and combination bands all occurring below 600 cm^{−1}, results in a very rich spectrum. It is known to be a major contributor to spectral line confusion in ground-based observations and is expected to complicate observations by Herschel, SOFIA and ALMA, making it imperative to fully characterize the entire spectrum. The lowest in-plane bend, ν_{13}, is 206.9(0.5) cm^{−1},^{b} and the first excited torsional state, ν_{21}, which is just 186 GHz above, have been detected in hot cores with antenna temperatures of a few Kelvin.^{c} The close proximity of ν_{13} and ν_{21}, as well as their low-lying nature, offers a unique opportunity to study the vibration-torsion-rotation coupling problem in the case of two nearly degenerate vibrational states. As expected from C_{s} symmetry and their A′ and A′′ nature, these states exhibit strong a− and b−symmetry Coriolis interactions, as well as interactions resulting from different sets of Eckhart-Sayvetz conditions being required in ν_{13} and ν_{21}. In the present work, the ν_{13} and ν_{21} states of propionitrile have been analyzed to high frequency and angular momentum quantum number. The spectrum, molecular constants,and insights into the vibration-torsion-rotation problem will be discussed.

aResearch described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contracts and cooperative agreements with the National Aeronautics and Space Administration.

aH. M. Heise, H. Lutz & H. Dreizler,Z.Nat.,29a,1345 (1974).

bH. M. Heise, F. Winther & H. Lutz,J. Mol. Spectrosc.,90,531 (1981).

cD. M. Mehringer, J. C. Pearson, J. Keene & T. G. Phillips,Ap.J.,608,306 (2004).